# Learning Prompt-Enhanced Context Features for Weakly-Supervised Video Anomaly Detection

### v1

@@ III-B Temporal Context Aggregation Module
The TCA module aims to capture both global and local dependencies of snippets to model the complete temporal relationship, as shown in Figure 4. Following the self-attention mechanism [17], the snippet feature $\mathbf{X}$ is first projected to the latent space by different linear layers, and the similarity matrix $\mathbf{M}$ is obtained by computing the inner product as follows:
@@
where $f_{q}(\cdot)$, $f_{k}(\cdot)$ and $f_{v}(\cdot)$ are three different linear mapping layers, $\top$ refers to the transpose operation, and $D_{h}$ is the hidden dimension in the latent space. Softmax normalization is then applied to generate the global attention map $\mathbf{A}^{g}$. The projected snippet feature is re-weighted by the attention map to obtain the global context feature $\mathbf{X}^{g}$.
Figure 4: Structure of the TCA module. DPE denotes dynamic position encoding.
Although the above operation facilitates global context modeling, it inevitably introduces long-range noise. To address this issue, we implement local context calibration by reusing the similarity matrix with a masking window. The process is denoted as:
@@
where $w$ is the window size of the mask and $T$ is the maximum length of the input sequence. Similarly, the projected snippet feature is re-weighted with the attention map $\mathbf{A}^{l}$ to obtain local calibration features $\mathbf{X}^{l}$, which can effectively capture slight changes and achieve feature enhancement in the local neighborhood.
Subsequently, a learnable weight is used to achieve local-global adaptive fusion, which is formulated as follows:
@@
where $\alpha$ and $1-\alpha$ denote the global and local weight for context fusion, respectively, and $\text{Norm}(\cdot)$ denote a combination of power normalization [50] and L2 normalization. Then a linear layer $f_{h}(\cdot)$ followed by residual connection and layer normalization is applied to obtain context feature $\mathbf{X}^{c}$.
@@
where $i$ and $j$ denote the absolute positions of two snippets, and $\gamma$ and $\beta$ are learnable weights and bias terms. In particular, the DPE is embedded into the similarity matrix $\mathbf{M}$ as a location prior, i.e., $\mathbf{M}\leftarrow\mathbf{M}+\mathbf{G}$, thus avoiding affecting the original feature distribution. Since DPE does not require a preset maximum sequence length, it is more flexible for variable-length videos. Also, it can suppress long-distance noises due to the properties of Gaussian-like kernel function.
III-C Multilayer Perceptron and Causal Convolution
To obtain high-level semantic representations, a two-layer MLP is utilized for feature reduction, as shown in Figure 5. Each Conv1D layer is followed by a GELU activation and dropout operation. This process is denoted as follows:
@@ III-C Multilayer Perceptron and Causal Convolution
Figure 5: Structure of MLP, CC module (left), and MIL-based loss (right).
@@ III-D Prompt-Enhanced Learning
The PEL module aims to enrich the visual representations by incorporating knowledge-based contextual information, thereby enhancing the model’s ability to identify anomalies in intricate scenarios. Specifically, the PEL module comprises three steps, namely prompt construction, fore-background separation, and cross-modal alignment, depicted in Figure 6.
@@ III-D1 Prompt Construction
Considering the versatility of prompts, we first select 12 common relations from ConceptNet as the pre-retrieval semantics, followed by choosing the top five items with the highest occurrence frequency across all categories as the retrieval relations, denoted as $\{r_{j}\}_{j=1}^{R}$. The concept dictionary $\mathcal{D}$ is then constructed by retrieving all edges of the relation $r_{j}$ established with a given class $c$ as the head or tail node. The non-category node in each edge is used as the key, and the relevance score of the edge is taken as the value. To remove noisy entries that may interfere with the semantics of anomalies, we first eliminate concepts with relevance scores less than or equal to 0 (Step 1). Next, the remaining entries are filtered by using the average of the relevance scores as the threshold (Step 2). In general, the higher the relevance score, the stronger the semantic relationship between the nodes.
@@
III-D2 Fore-Background Separation
Since snippet-level features lack full contextual information, direct alignment with prompt features may result in boundary deviations of anomaly localization. Therefore, it is crucial to distinguish between class-specific foreground and class-agnostic background from consecutive snippets. To achieve this, we utilize the scaled anomaly scores as activations to produce video-level foreground and background features, respectively, represented as follows:
@@ III-D2 Fore-Background Separation
Figure 6: Illustration of PEL module. knowledge-based prompts are first constructed via ConceptNet. Video-level foreground and background are then separated. Cross-modal alignment encourages video-prompt pairs with the same semantics to converge and form discriminative decision boundaries.
@@ III-D3 Cross-modal Alignment
Finally, we present our approach to enhancing the fine-grained semantics of visual features by aligning them with the prompt features, as illustrated in Figure 7. For the abnormal video, the visual foreground $\mathbf{V}^{fg}_{a}$ is aligned to its knowledge-based prompt feature $\mathbf{T}_{a}^{c}$, while the background $\mathbf{V}^{bg}_{a}$ is aligned to the unified normal prompt $\mathbf{T}^{c}_{n}$. For normal video, only the alignment between the normal foreground $\mathbf{V}^{fg}_{n}$ and the corresponding prompt $\mathbf{T}^{c}_{n}$ is considered. The process is formulated as follows:
@@
where $\mathbf{V}\in\left\{\mathbf{V}_{a}^{fg}\bigcup\mathbf{V}_{a}^{bg}\bigcup\mathbf{V}_{n}^{fg}\right\}$ and $\mathbf{T}\in\left\{\mathbf{T}_{a}^{c}\bigcup\mathbf{T}_{n}^{c}\right\}$ denote visual and textual representation, respectively, $C+1$ denotes $C$ abnormal classes and 1 normal class, and $\tau$ is a temperature coefficient. The cross-modal alignment loss is computed using the Kullback-Leibler divergence, as follows:
@@
III-E Objective Function and Score Smoothing
@@
During the testing phase, we introduce the SS module to reduce false alarms caused by frame jitter or switching. This module employs distinct pooling strategies to achieve optimal results. Given an anomaly score sequence $\{s_{i}\}_{i=1}^{T}$ and a pooling window of size $\kappa$, the smoothing process can be expressed as follows:
@@ IV-B2 Hyperparameter settings
The hidden dimension of the TCA module is set to 128, and the fusion weight $\alpha$ is initialized to 0.5. The two Conv1D layers of MLP have 512 and 300 nodes, respectively, both with a dropout rate of 0.1. For the UCF-Crime, XD-Violence, and ShanghaiTech datasets, we adopt local window sizes of 9, 9, and 5, respectively. The causal convolution kernel size $\Delta t$ is set to 9, 3 and 3, respectively, while the temperature coefficient $\tau$ is initialized to 0.09, 0.05, and 0.2, respectively. The scaling factor $\mu$ for fore-background separation is set to 10, and the loss coefficient $\lambda$ is set to 1 for the UCF-Crime and XD-Violence datasets, while it is set to 9 for the ShanghaiTech dataset.
@@ IV-B3 Training and Test Details
During the training phase, our model is trained using the Adam optimizer [52] with a batch size of 128 and 50 epochs in total. The initial learning rate for the three datasets is set to $5\times 10^{-4}$, with a cosine decay strategy. For the balance of computational efficiency and detection performance, the snippet sampling threshold in the training phase is set to 200 as in [18] [12] [45] [53]. In the testing phase, we apply sliding pooling with window sizes of 7 and 3 to the UCF-Crime and ShanghaiTech datasets, respectively. Moving pooling with a window size of 9 is adopted for the XD-Violence dataset. More details can be found in supplementary materials.
@@ IV-C Comparison With State-of-the-Art Methods
Zhang et al. [15] I3D+VGGish 81.43 -
@@
Our model achieves competitive results across three datasets among all weakly supervised methods. Although our model’s AUC value on the UCF-Crime dataset is lower than that of UR-DMU [21] by 0.21%, we outperform UR-DMU on the XD-Violence dataset with an absolute gain of 3.93% AP and lower false alarm rates on both datasets. UR-DMU employs an additional encoder layer to learn local features with temporal masks, whereas our TCA module achieves local-global context modeling by reusing the similarity matrix. Additionally, UR-DMU introduces a dual memory unit to store normal and abnormal patterns and enlarges the decision boundary through a magnitude distance loss. In contrast, our proposed PEL module enhances fine-grained detection performance by considering the intra-class discriminability of exceptions while ensuring inter-class separability. Notably, our model even outperforms multimodal-based approaches, namely CMA-LA [60] and MACIL-SD [61], with 2.19% and 2.05% gains on the XD-violence dataset, respectively. Although the PEL module requires extra textual information in the training phase, our model only requires appearance features as input during inference. Furthermore, our model achieves a new state-of-the-art performance on the ShanghaiTech dataset, with a 0.66% improvement in AUC compared to the S3R [14] and CRFD [12] methods. For the first time, the false alarm rate is also reduced to 0, indicating that the model has sufficient confidence and robustness for normal samples.
@@ IV-D1 Contribution of Each Component
Table IV presents the contributions of each component, where ‘Baseline’ refers to the combination of MLP and CC module. Significantly, the TCA module is responsible for the primary improvement, delivering AUC gains of 9.77% and 5.57% on the ShanghaiTech and XD-Violence datasets, respectively, highlighting the need for context modeling. Furthermore, the introduction of the PEL module enhances the detection performance of our model on all three datasets, with a substantial AP gain of 1.98% on XD-Violence. We contend that this is due to the internal distribution characteristics of the dataset, which has a wider range of acquisition sources than surveillance videos such as UCF-Crime and ShanghaiTech, resulting in diverse and complex video content. Finally, the SS module further improves the detection performance of our model by 0.4%, 0.33%, and 0.14% on the three benchmarks, respectively, demonstrating the effectiveness and generality of the module.
@@
Learnable
$\alpha$ $1-\alpha$ 85.72 83.28 97.82
@@ IV-D2 Contribution of Components of the TCA module
Table V demonstrates that local context modeling can effectively improve the detection performance of our baseline. However, global context modeling resulted in a significant degradation of performance on the UCF-Crime and ShanghaiTech datasets, while outperforming local context modeling on XD-Violence. We assert that global context modeling introduces long-range noise to fixed surveillance videos, leading to reduced discrimination among snippet features. In contrast, film and TV videos require a global receptive field to capture complex semantics across snippets. The adaptive fusion of local and global contexts significantly enhanced the model’s performance, indicating that each approach plays a complementary role to the other, which outperformed the baseline by 9.54% and 5.37% on the XD-Violence and ShanghaiTech datasets, respectively. Finally, the improvement from DPE indicates the importance of relative distance as a location prior.
(a) AUC results w.r.t. sub-classes on UCF-Crime.
(b) AP results w.r.t. sub-classes on XD-Violence.
TABLE VIII: Contribution of node filtering in prompt construction
Method
UCF-Crime
AUC (%)
XD-Violence
AP (%)
None 85.34 84.35
step 1 85.87 84.61
step 1 + step 2 (fixed threshold) 86.31 85.10
step 1 + step 2 (dynamic threshold) 86.36 85.26
@@
Next, we report the superiority of adaptive fusion in Table VI. Compared with fixed weights, the learned weights yield better results across all three datasets, which enables the dynamic adjustment of the fusion ratio of the local-global context to suit diverse datasets. Table VII compares the performance of different normalization methods. The combination of power normalization and L2 normalization achieves the best results on UCF-Crime and ShanghaiTech. However, we observe that either power normalization or L2 normalization leads to a significant drop in performance for XD-Violence. We argue that normalization may jeopardize the diversity distribution of this dataset since it contains videos from different sources and types. While normalization is necessary for surveillance videos with fixed lenses, we remove the normalization operation in adaptive fusion for XD-Violence.
IV-D3 Contribution of Components of the PEL
First, we investigate the effectiveness of threshold filtering in prompt construction, as presented in Table VIII. Specifically, ‘step 1’ denotes the removal of entries with relevance scores not greater than 0, while ‘step 2’ filters the remaining entries by a specific threshold. The results reveal that ‘step 1’ improves the detection performance by 0.53% and 0.26% on the UCF-Crime and XD-Violence datasets, respectively, compared with that before filtering. Moreover, ‘step 2’ with dynamic threshold further enhances the performance on both datasets, indicating the successful removal of redundant entries. The dynamic threshold ensures the completeness of relevant semantic relationships to a certain extent while the fixed threshold may introduce low-relevance nodes with potentially anomalous semantics.
We also compare different prompt templates in Table IX. Using category labels as prompt templates improves the detection performance by 0.23% and 1.41% on the UCF-Crime and XD-Violence datasets, respectively, outperforming the hand-crafted templates (the $\text{2}^{\text{nd}}$ and $\text{3}^{\text{rd}}$ rows) and WordNet definitions (the $\text{4}^{\text{th}}$ row). However, our ConceptNet-based prompts outperform the label-only prompts by 0.41% and 0.57% on the two benchmarks, respectively. Retrieving class-specific concepts based on semantic relations facilitates the mining of potential co-occurring anomalous items, while threshold filtering ensures strongly related semantic concepts, aiding the capture of fine-grained visual information.
TABLE IX: performance comparison of different prompt template
Prompt Template
UCF-Crime
AUC (%)
XD-Violence
AP (%)
{label} 85.95 84.69
{‘a video of’ + label} 85.55 84.51
{‘a long video of’ + label} 85.63 84.59
{label + WordNet Definition} [47] 85.92 84.14
{Learnable Prompt + label} [26] 85.48 84.48
{label + ConceptNet Relation} 86.36 85.26
TABLE X: contribution of fore-background separation in the PEL module
Method
UCF-Crime
AUC (%)
XD-Violence
AP (%)
w/o F-B separation 85.71 79.79
w/ F-B separation 86.36 85.26
TABLE XI: contribution of score smoothing module
Method
UCF
AUC (%)
XD
AP (%)
SHTech
AUC (%)
None 86.36 85.26 98
moving window 86.65 85.59 98.03
sliding window 86.76 85.56 98.14
Finally, we investigate the contribution of foreground-background separation, as presented in Table X. ‘w/o F-B separation’ denotes the mean pooling operation of snippet features and alignment with the corresponding prompt features, which leads to a significant performance degradation. The mean pooling operation weakens the anomalous region and introduces extra background noise, resulting in a biased cross-modal alignment. Foreground-background separation effectively suppresses background interference, ensuring semantic consistency between visual foreground and prompt representations.
IV-D4 Contribution of PEL to Fine-grained Anomaly Detection
We report the detection results of different anomaly subclasses after cross-modal alignment in Figure 7. For the UCF-Crime dataset, over half of the categories demonstrated improvements in AUC, with ‘explosion’ gaining 8.5% and ‘vandalism’ gaining 5.19%. However, some subclasses experienced significant drops in performance, with ‘abuse’ decreasing by 9.96% and ‘shooting’ by 2.71%. We discovered that ConceptNet’s relevant entries include ‘disrespect’, ‘vituperation’, and ‘scurrility’, while the term ‘abuse’ in UCF-Crime refers to the mistreatment of animals by humans. We contend that the polysemy of English words results in a deviation of the visual representation from the original context after cross-modal alignment. Overall, the macro-AUC of the 13 subclasses of the UCF-Crime dataset increased by 1.05%.
@@ IV-D4 Contribution of PEL to Fine-grained Anomaly Detection
For the XD-Violence dataset, all subclasses except for ‘car accident’ demonstrated varying degrees of increase in AP. Interestingly, the AP of ‘abuse’ increased significantly by 9.04%, which is the opposite of the UCF-Crime dataset. This is due to the former being richer in video types, including verbal and physical conflicts between characters, which is consistent with the concept of nodes extracted by ConceptNet. Furthermore, the mAP of all 6 anomaly subclasses improved by 2.17%, further supporting the contribution of cross-modal alignment to fine-grained anomaly detection.
IV-D5 Contribution of SS
Table XI illustrates the impact of various pooling windows in the SS module. The utilization of both moving and sliding pooling techniques effectively enhances the model’s performance while simultaneously reducing the false alarm rate. The ShanghaiTech and UCF-Crime datasets exhibit favorable outcomes with the latter, whereas the former yields superior results for the XD-Violence dataset.
IV-E Analysis of Time–Space Complexity
In this subsection, we compared the performance and efficiency with recent temporal modeling methods, as shown in Table XII. To ensure a fair comparison, we reproduce existing modules using open-source code and embed them in our baseline. Specifically, HL-Net [18] models global and local relationships separately using two parallel graph convolutional networks and concatenates the features for fusion. MTCM [45] employs multiple 1D convolution layers to capture multi-scale context information, while RTFM [13] captures multi-level temporal relations using a self-attention mechanism with a pyramid dilated convolution block. However, these multi-branch structures unavoidably introduce additional parameters and computation costs. Despite maintaining the same space-time complexity as the LGTE [65], our TCA module outperforms the former by 1.38% and 4.21% on the UCF-Crime and XD-Violence datasets, respectively. We argue that although LGTE uses channel grouping to improve encoding efficiency, it may compromise the complete semantics of feature channels. In contrast, the TCA module preserves contextual continuity by reusing the similarity matrix, and adaptive aggregation and dynamic positional coding further enhance the discriminative nature of snippet features. Notably, the feature extraction model I3D can achieve a processing speed of 263 FPS on a single Tesla A40 GPU when the input frame resolution is set to $224\times 224$. Subsequently, the anomaly detection model outputs anomaly scores at 350 FPS (2.9 ms), striking a good balance between detection performance and inference efficiency.
TABLE XII: Comparison of the spatio-temporal complexity of different methods

### v2

@@ III-B Temporal Context Aggregation Module
As previously discussed, the Temporal Context Aggregation (TCA) module, depicted in Figure 2, diverges from the conventional multi-branch parallel frameworks used in existing methods. It concurrently models global-local dependencies by leveraging a similarity matrix and adaptive fusion. The snippet feature $\mathbf{X}$ is first projected to the latent space by different linear layers, and the similarity matrix $\mathbf{M}$ is obtained by computing the inner product as follows:
@@
where $f_{q}(\cdot)$, $f_{k}(\cdot)$ and $f_{v}(\cdot)$ are three different linear layers, $\top$ refers to the transpose operation, and $D_{h}$ is the hidden dimension in the latent space. Softmax normalization is then applied to generate the global attention map $\mathbf{A}^{g}$. The projected snippet feature is re-weighted by the attention map to obtain the global context feature $\mathbf{X}^{g}$.
Although the above operation facilitates global context modeling, it inevitably introduces long-range noise. To address this issue, we implement local context calibration by reusing the similarity matrix with a masking window, which can be formulated as:
@@
where $w$ is the window size of the mask and $T$ is the maximum length of the input sequence. Eq. (4) ensures that the $i^{th}$ snippet interacts only with its neighborhood of window size $w$. The lower bound of this window is the earliest moment that can be historically observed, and the upper bound is the maximum length of the sequence. Similarly, the projected snippet feature is re-weighted with the attention map $\mathbf{A}^{l}$ to obtain local calibration features $\mathbf{X}^{l}$, which can effectively capture slight changes and achieve feature enhancement in the local neighborhood.
Subsequently, we resort to a learnable fashion rather than direct concatenation or average pooling to achieve local-global contextual adaptive fusion, allowing the model to dynamically balance the importance of global temporal patterns and local nuances. The fusion process is formulated as follows:
@@
where $\alpha$ and $1-\alpha$ denote the global and local weight for context fusion, respectively, and $\text{Norm}(\cdot)$ denote a combination of power normalization [52] and L2 normalization. Then a linear layer $f_{h}(\cdot)$ followed by residual connection and layer normalization $\text{LN}(\cdot)$ is applied to obtain context feature $\mathbf{X}^{c}$.
@@
where $i$ and $j$ denote the absolute positions of two snippets, and $\gamma$ and $\beta$ are learnable weights and bias terms. In particular, the DPE is embedded into the similarity matrix $\mathbf{M}$ as a location prior, i.e., $\mathbf{M}\leftarrow\mathbf{M}+\mathbf{G}$, thus avoiding affecting the original feature distribution. Unlike fixed position encodings in [17], DPE adapts to varying video lengths due to its dynamic nature. Also, the Gaussian-like kernel of DPE inherently suppresses the influence of long-distance noise, which emphasizes closer snippet relationships over distant ones implies an innate sensitivity to non-linear patterns.
III-C Multilayer Perceptron and Classifier
To obtain high-level semantic representations, a two-layer MLP is utilized for feature reduction. Each Conv1D layer is followed by a GELU activation and dropout operation. This process is denoted as follows:
@@ III-D Prompt-Enhanced Learning
The PEL module aims to enrich the visual representations by incorporating knowledge-based contextual information, thereby enhancing the model’s ability to identify anomalies in intricate scenarios. Specifically, the PEL module comprises three steps, namely prompt construction, context separation, and cross-modal alignment, as depicted in Figure 2.
@@ III-D1 Prompt Construction
[…]
@@
Figure 3: An example of concept dictionary given the anomaly class fighting. The arrows point from the head node to the tail node with relevance scores, and the colors indicate different relationships. The entire graph constitutes a concept dictionary, where the bold items are those retained after node filtering.
III-D2 Context Separation
Since snippet-level features don’t encapsulate complete contextual details, aligning them directly with prompt features could lead to inaccuracies in anomaly localization. Therefore, it’s imperative to differentiate between class-specific foreground and class-agnostic background in successive snippets. We employ scaled anomaly scores as activations to generate video-level foreground and background features. These are represented as follows:
@@ III-D3 Cross-modal Alignment
Finally, we present our approach to enhancing the fine-grained semantics of visual features by aligning them with the prompt features, as illustrated in Figure 2. For abnormal videos, foreground features $\mathbf{V}^{fg}_{a}$ are matched with the corresponding abnormal prompt $\mathbf{T}_{a}^{c}$, increasing the chances of accurate abnormal behavior identification. In contrast, background features $\mathbf{V}^{bg}_{a}$ are paired with the normal prompt $\mathbf{T}^{c}_{n}$, maintaining normal classification within an abnormal setting. For semantically inconsistent video-prompt pairs, repulsion is achieved by minimizing the cosine distance of all negative pairs, further forming discriminative decision boundaries. The process is formulated as:
@@
where $\psi(\cdot)$ measures the cosine similarity between the visual representation $\mathbf{V}\in\left\{\mathbf{V}_{a}^{fg}\bigcup\mathbf{V}_{a}^{bg}\bigcup\mathbf{V}_{n}^{fg}\right\}$ and the textual representation $\mathbf{T}\in\left\{\mathbf{T}_{a}^{c}\bigcup\mathbf{T}_{n}^{c}\right\}$. Probability $p^{v2t}(\cdot)$ estimates how likely a visual feature matches a specific prompt across $C$ anomaly classes and 1 normal class, where $\tau$ is a temperature coefficient. Finally, the cross-modal alignment loss is computed using the Kullback-Leibler divergence, which forces the network to learn to distinguish between the visual content of the video that represents abnormal behavior (foreground) and the content that is not relevant to the abnormal behavior (background). The loss function is formulated as follows:
@@
III-E Training and testing procedures
@@
Considering the temporal consistency, we utilize a score smoothing (SS) strategy in the testing phase to mitigate the impact of transient noise that does not conform to the expected duration of an anomaly. This strategy employs distinct pooling operations to achieve optimal results. Given an anomaly score sequence $\{s_{i}\}_{i=1}^{T}$ and a pooling window of size $\kappa$, the smoothing process can be expressed as follows:
@@ IV-B2 Hyperparameter settings
The hidden dimension of the TCA module is set to 128, and the fusion weight $\alpha$ is initialized to 0.5. The two Conv1D layers of MLP have 512 and 300 nodes, respectively, both with a dropout rate of 0.1. For the UCF-Crime, XD-Violence, and ShanghaiTech datasets, we adopt local window sizes of 9, 9, and 5, respectively. The causal convolution kernel size $\Delta t$ is set to 9, 3 and 3, respectively, while the temperature coefficient $\tau$ is initialized to 0.09, 0.05, and 0.2, respectively. The scaling factor $\mu$ for context separation is set to 10, and the loss coefficient $\lambda$ is set to 1 for the UCF-Crime and XD-Violence datasets, while it is set to 9 for the ShanghaiTech dataset.
@@ IV-B3 Training and Test Details
During the training phase, our model is trained using the Adam optimizer [54] with a batch size of 128 and 50 epochs in total. The initial learning rate for the three datasets is set to $5\times 10^{-4}$, with a cosine decay strategy. For the balance of computational efficiency and detection performance, the snippet sampling threshold in the training phase is set to 200 as in [18] [12] [47] [55]. In the testing phase, we apply sliding pooling with window sizes of 7 and 3 to the UCF-Crime and ShanghaiTech datasets, respectively. Moving pooling with a window size of 9 is adopted for the XD-Violence dataset. More details can be found in our supplementary materials.
@@ IV-C Comparison With State-of-the-Art Methods
Cho et al.[40] I3D RGB 81.30 -
ACF [61] I3D+VGGish 80.13 -
MSAF [45] I3D+VGGish 80.51 -
CUPL [15] I3D+VGGish 81.43 -
@@
Our model achieves competitive results across three datasets among all weakly supervised methods. Although our model’s AUC value on the UCF-Crime dataset is lower than that of UR-DMU [21] by 0.21%, we outperform UR-DMU on the XD-Violence dataset with an absolute gain of 3.93% AP and lower false alarm rates on both datasets. UR-DMU employs an additional encoder layer to learn local features with temporal masks, whereas our TCA module achieves local-global context modeling by reusing the similarity matrix. Additionally, UR-DMU introduces a dual memory unit to store normal and abnormal patterns and enlarges the decision boundary through a magnitude distance loss. In contrast, our proposed PEL module enhances fine-grained detection performance by considering the intra-class discriminability of exceptions while ensuring inter-class separability. Notably, our model even outperforms multimodal-based approaches, namely CMA-LA [46] and MACIL-SD [44], with 2.19% and 2.05% gains on the XD-violence dataset, respectively. Although the PEL module requires extra textual information in the training phase, our model only requires appearance features as input during inference. Furthermore, our model achieves a new state-of-the-art performance on the ShanghaiTech dataset, with a 0.54% improvement in AUC compared to Cho et al [40]. For the first time, the false alarm rate is also reduced to 0, indicating that the model has sufficient confidence and robustness for normal samples.
@@ IV-D1 Contribution of Each Component
Table IV details the impact of each model component. The ‘Baseline’, comprising MLP and the Classifier, is notably enhanced by the TCA module, which yields substantial improvements of 9.77% and 5.57% on XD-Violence and ShanghaiTech datasets, respectively. This underscores the importance of context modeling. The PEL module further advances detection across all datasets, particularly notable with a 1.98% AP increase on XD-Violence. This reflects the varied content sources of XD-Violence compared to UCF-Crime and ShanghaiTech. The SS strategy also demonstrates consistent, albeit smaller, improvements across the benchmarks, confirming its effectiveness and versatility.
@@
Learnable $\alpha$ $1-\alpha$ 85.72 83.28 97.82
@@ IV-D2 Contribution of Components of the TCA module
Table V shows that local context modeling significantly improves baseline detection. In contrast, global context modeling decreases performance on UCF-Crime and ShanghaiTech but is more effective on XD-Violence. We argue that global context may introduce noise in fixed surveillance footage, reducing feature discrimination, but is beneficial for diverse film and TV content. Combining local and global contexts outperforms the baseline by 9.54% and 5.37% on XD-Violence and ShanghaiTech, respectively. This indicates the complementary roles of these approaches. The improvement from DPE highlights the importance of relative distance in feature location.
@@
Furthermore, Table VI demonstrates adaptive fusion’s superiority over fixed weights, showing better results across all datasets. This supports the dynamic adjustment of local-global context fusion to suit diverse videos. Table VII compares normalization methods. Power and L2 normalization combined yield optimal results on UCF-Crime and ShanghaiTech. However, on XD-Violence, either normalization method significantly reduces performance, likely affecting the dataset’s diverse video source distribution. Consequently, we omit normalization for XD-Violence in adaptive fusion to preserve content diversity.
Finally, we visualize the correlation of context features to demonstrate the encoding performance of the TCA module. As shown in Figure 4, the heatmap demonstrates the cosine similarity between the snippets. From this, we observe: 1) UCF-Crime videos show less variability compared to XD-Violence due to their fixed surveillance nature versus XD-Violence’s complex scene and camera movements; 2) Local features offer more discriminability than global features, which tend to introduce more long-range noise; and 3) The aggregation of local and global features suppresses redundancies and enhances snippet discriminability.
@@
TABLE VIII: Comparison of the spatio-temporal complexity of different methods
@@
IV-D3 Analysis of Time–Space Complexity
This subsection evaluates the Temporal Context Aggregation (TCA) module’s performance and efficiency against recent temporal modeling techniques, shown in Table VIII. For a fair comparison, we used open-source codes to reproduce existing modules and integrated them into our baseline. HL-Net [18] uses dual graph convolutional networks for global and local relationship modeling. MTCM [47] employs multiple 1D convolution layers for multi-scale context, and RTFM [13] uses self-attention with pyramid dilated convolution for understanding temporal relations. These methods significantly increase parameters and computational demands. However, our TCA module, with similar space-time complexity, outperforms LGTE [65] by 1.38% and 4.21% on UCF-Crime and XD-Violence datasets, respectively. The TCA module’s success is due to its effective use of the similarity matrix, maintaining contextual continuity. This contrasts with LGTE’s channel grouping, which might lose some feature channel details. TCA’s adaptive aggregation and dynamic positional coding also improve snippet feature uniqueness. Notably, the feature extraction model I3D can achieve a processing speed of 263 FPS on a single Tesla A40 GPU when the input frame resolution is set to $224\times 224$. Subsequently, the anomaly detection model outputs anomaly scores at 350 FPS (2.9 ms per frame), striking a good balance between detection performance and efficiency.
TABLE IX: Contribution of node filtering in prompt construction
Method
UCF-Crime
AUC (%)
XD-Violence
AP (%)
None 85.34 84.35
step 1 85.87 84.61
step 1 + step 2 (fixed threshold) 86.31 85.10
step 1 + step 2 (dynamic threshold) 86.36 85.26
TABLE X: performance comparison of different prompt template
Prompt Template
UCF-Crime
AUC (%)
XD-Violence
AP (%)
{label} 85.95 84.69
{‘a video of’ + label} 85.55 84.51
{‘a long video of’ + label} 85.63 84.59
{label + WordNet Definition} [49] 85.92 84.14
{Learnable Prompt + label} [26] 85.48 84.48
{label + ConceptNet Relation} 86.36 85.26
IV-D4 Contribution of Components of the PEL
Table IX first evaluates the impact of node filtering in prompt construction. Step 1 involves removing entries with relevance scores at or below zero, while Step 2 applies a specific threshold to the remaining entries. Step 1 enhances detection by 0.53% and 0.26% on UCF-Crime and XD-Violence datasets, respectively. Step 2 with a dynamic threshold further improves performance on both datasets, indicating effective reduction of redundant entries. The dynamic threshold maintains relevant semantic relationships, while a fixed threshold might include low-relevance, potentially anomalous entries.
Prompt Templates Comparison
Different prompt templates are compared in Table X. Category labels as prompts increase detection by 0.23% and 1.41% on UCF-Crime and XD-Violence, respectively, surpassing hand-crafted templates and WordNet definitions. ConceptNet-based prompts further improve performance by 0.41% and 0.57% on these datasets. Using class-specific concepts based on semantic relations aids in identifying co-occurring anomalous items. Threshold filtering ensures the selection of highly relevant semantic concepts, enhancing fine-grained visual information capture.
TABLE XI: contribution of context separation in the PEL module
Method
UCF-Crime
AUC (%)
XD-Violence
AP (%)
w/o context separation 85.71 79.79
w/ context separation 86.36 85.26
TABLE XII: contribution of score smoothing strategy
Method
UCF
AUC (%)
XD
AP (%)
SHTech
AUC (%)
None 86.36 85.26 98
moving window 86.65 85.59 98.03
sliding window 86.76 85.56 98.14
Context Separation
As shown in Table XI, ‘w/o context separation’ denotes the mean pooling operation of snippet features and alignment with the corresponding prompt features, leading to notable performance drops. Mean pooling dilutes the anomalous region’s impact and introduces background noise, skewing cross-modal alignment. Therefore, context separation effectively minimizes background noise, ensuring better semantic consistency between visual foreground and prompt representations.
IV-D5 PEL’s Impact on Fine-Grained Anomaly Detection
As shown in Figure 5(a), our PEL method demonstrates superior anomaly detection capabilities on the UCF-Crime dataset, outperforming RTFM [13] and UMIL [41] in most categories with significant leads in Abuse (76.9%) and Assault (96.2%). Its robust average AUC of 72.2% underscores its effectiveness and potential applicability in security and surveillance. While PEL excels in complex anomaly scenarios, it lags slightly in Explosion, suggesting a need for refinement in detecting abrupt anomalies. In Figure 5(b), the performance of the PEL method in detecting various sub-classes of violence is contrasted with a baseline (w/o PEL). PEL consistently improves detection across all categories, with significant AP gains in Fighting (83.8% vs. 81.0%) and Abuse (70.5% vs. 61.5%). These improvements suggest that PEL effectively captures the nuances of complex, dynamic events. The overall average performance increase to 70.3% from 68.1% also underscores PEL’s robustness across diverse scenarios. However, in Car Accident, the impact of PEL is minimal, indicating potential areas for refinement.
IV-D6 Impact of Score Smoothing
Table XII illustrates the effects of different pooling window sizes in the SS (Sliding and Sliding) strategy. Implementing both moving and sliding pooling techniques has shown to not only improve the model’s performance but also decrease the false alarm rate. Specifically, the sliding pooling technique yields better results for ShanghaiTech and UCF-Crime datasets. Conversely, the moving pooling technique is more effective for the XD-Violence dataset.
(a) Class-wise AUC results of three methods on UCF-Crime.
(b) Class-wise AP results of before/after PEL on XD-Violence.
Figure 6: Model performance under different hyperparameter settings. (a) and (b) show the AUC and FAR results of our model for different local windows $w$ and causal convolution size $\Delta t$, respectively. (c) and (d) illustrate the effects of two different score smoothing methods on the model’s performance as the window size varies.